OpenStudy (anonymous):

What is the radius of (x - 3)2 + (y + 4)2 = 14?

OpenStudy (anonymous):

mathway.com

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] has center \((h,k)\) and radius \(r\) in your case \(r^2=14\)

OpenStudy (nirmalnema):

The general form of the equation for a circle with its center at (h, k) and radius of r is: \[(x-h)^2 +(y-k)^2=r^2\]

OpenStudy (nirmalnema):

r^2=14 r=\[\sqrt{14}\]